A thin film of acetone (n = 1.25) coats a thick glass plate (n = 1.50). White light is incident normal to the film. In the reflections, fully destructive interference occurs at 560 nm and fully constructive interference at 630 nm. Calculate the thickness of the acetone film.
Answer #1

For constructive interference,

2 nt = m lambda

t = m x 630/(2 x 1.25) = 252 m ...... (1)

For destructive interference,

t = (m + 0.5) x 560/(2 x 1.25) = 224 (m + 0.5) ..... (2)

From (1) and (2)

224 m + 112 = 252 m

m = 4

So, t = 252 x 4 = 1008 nm

Comment in case any doubt please rate my answer...